OpenStudy (anonymous):

x^2+6x-7=0 solve for x

OpenStudy (anonymous):

(x+7)(x-1)=0............x=-7,+1

OpenStudy (anonymous):

x=-1, x=7

OpenStudy (anonymous):

x^2+6x-7=0 <==> (x+7)(x-1)=0 ==> x=-7 or x=1